OpenStudy (anonymous):

3. Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete senten

OpenStudy (anonymous):

french train 89.44meters/second there are 60*60 seconds in an hour 89.44*60*60=321984meters/hour divide by 1000 to change to kilometers ....321.984km/hr France =321.98km/hr japan=320km/hr ellen =321 km/hr to be second fastest.. you are asked for miles/hr 321*0.621=199.34 miles/hr as..1km=0.621 miles distance =speed*time ....199.34*2.5=498.35 miles in 2.5 hrs